# Bias Detection via Maximum Subgroup Discrepancy

### v1

@@ 2.4 Mixed-integer optimization
Mixed-integer optimization (MIO, [55]) is a powerful framework for modeling and solving mathematical optimization problems, where some decision variables take values from a discrete set while others are continuously valued. Despite MIO being a general framework for solving NP-hard problems, the MIO solvers are speeding up by approximately 22% every year, excluding hardware speedups [30]. We use the abbreviation MIO, though we consider only mixed-integer linear formulations.
@@
MIO is used in machine learning, especially when one optimizes over discrete measures or decisions. This includes learning logical rules, including DNFs. Malioutov and Varshney [35] learn DNFs through a sequential generation of terms using an LP relaxation of an MIO formulation. Later, Wang and Rudin [51] used MIO to optimize full DNFs, and Su et al. [45] introduced a formulation with Hamming distance as an alternative objective function. A crucial improvement to the scalability of exact learning of DNFs was the BRCG [10] utilizing column generation to generate candidate terms. Recently, MIO was utilized to learn a DNF classifier with fairness constraints [32]. Importantly, we optimize only a single term (conjunction), representing the maximally discrepant subgroup, to evaluate bias. We do not perform predictions.
@@
As an aside, note that [38] (see also [21]) use DNFs to compare two models. This is done by building two DNFs as proxy models and comparing them. However, we could also directly compare the distributions of the two models, as proposed here.
3 Maximum Subgroup Discrepancy
@@
Let $\mathcal{X}\subset\mathbb{R}^{d}$ be an input space with $d$ features and let $\mathcal{P}\subseteq[d]$ be a subset of features that are protected attributes. Throughout the rest of the paper, we assume that the protected attributes are binary, which can always be achieved by quantization and one-hot encodings. For a vector $x\in\mathbb{R}^{d}$ and $p\in\mathcal{P}$ let $x_{p}\in\left\{0,1\right\}$ denote its $p$-th coordinate, and let $\bar{x}_{p}:=1-x_{p}$ denote its logical negation. The set of functions $\mathcal{L}=\left\{x_{p},\bar{x}_{p}\right\}_{p\in\mathcal{P}}$ is called literals. A term is a conjunction of literals. That is, for a subset $S\subset\mathcal{L}$, the term $\chi_{S}$ is defined as a function $\mathbb{R}^{d}\rightarrow\left\{0,1\right\}$ given by
$\chi_{S}(x)=\prod_{s\in S}s(x).$ (3)
@@
The subsets $S\subset\mathcal{L}$ are called subgroups, and with some abuse of notations we also refer to the corresponding parts of the data defined by $S$,
$\mathcal{X}_{S}=\left\{x\in\mathbb{R}^{d}\hskip 5.69054pt|\hskip 5.69054pt\chi_{S}(x)=1\right\},$ (4)
@@
With this notation, the $\mathrm{MSD}_{\mathrm{diff}}$ distance on feature set $\mathcal{P}$ between two distributions, $\mu,\nu$, is defined as
$\mathrm{MSD}_{\mathrm{diff}}(\mu,\nu;\mathcal{P})=\sup_{S\in\mathcal{S}_{\mathcal{P}}}\left|\mu(\chi_{S})-\nu(\chi_{S})\right|.$ (5)
@@
$\mathcal{B}_{\mathcal{P}}=\left\{\prod_{p\in\mathcal{P}}l_{p}\hskip 5.69054pt|\hskip 5.69054pt\text{ where $l_{p}=x_{p}$ or $l_{p}=\bar{x}_{p}$}\right\}\subset\mathcal{S}_{\mathcal{P}}.$ (6)
@@
$\left\|\mu-\nu\right\|_{\infty}=\sup_{S\in\mathcal{B}}\left|\mu(\chi_{S})-\nu(\chi_{S})\right|.$ (7)
@@
Let $\mathcal{F}_{\mathcal{P},\infty}$ be the set of functions $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ that depend only on the coordinates in $\mathcal{P}$, and are bounded by $1$. Then, the total variation222marginalised to $\mathcal{P}$ (TV), may be written as
$\left\|\mu-\nu\right\|_{TV}$ $=\frac{1}{2}\sup_{f\in\mathcal{F}_{\mathcal{P},\infty}}\left|\mu(f)-\nu(f)\right|$ (8)
$=\frac{1}{2}\sum_{S\in\mathcal{B}}\left|\mu(\chi_{S})-\nu(\chi_{S})\right|,$ (9)
@@
$\left\|\mu-\nu\right\|_{\infty}\leq\mathrm{MSD}_{\mathrm{diff}}(\mu,\nu;\mathcal{P})\leq\left\|\mu-\nu\right\|_{TV},$ (10)
@@
We thus observe that $\mathrm{MSD}_{\mathrm{diff}}(\mu,\nu;\mathcal{P})$ is a stronger distance than $\ell_{\infty}$, which allows us to consider subgroups with partially specified values. At the same time, it is not as strong as the total variation, which requires the sum over all base term differences to be small (rather than each of the differences being small individually). On the other hand, as discussed in Section 1, TV is, in fact, too strong to be practically computable due to its high sample complexity, while $\mathrm{MSD}_{\mathrm{diff}}$ has a sample complexity linear in the number of attributes.
3.1 Sample Complexity
In this section, we prove Theorem 3.1, which allows us to quantify the error of estimating $\mathrm{MSD}_{\mathrm{diff}}(\mu,\nu;\mathcal{P})$ from finite samples of these distributions.
@@ 3.1 Sample Complexity
$\hat{\mu}=\frac{1}{N_{1}}\sum_{i}\delta_{x_{i}}\text{ and }\hat{\nu}=\frac{1}{N_{2}}\sum_{i}\delta_{x^{\prime}_{i}},$ (11)
@@
$\mathrm{MSD}_{\mathrm{diff}}(\mu,\nu;\mathcal{P})\leq\mathrm{MSD}_{\mathrm{diff}}(\hat{\mu},\hat{\nu};\mathcal{P})+4\sqrt{\frac{2\left|\mathcal{P}\right|+\log\frac{2}{\delta}}{2N}}.$ (12)
In words, if the number of samples $N$ is of the order of the number of protected attributes, $\left|\mathcal{P}\right|$, or larger, we can estimate $\mathrm{MSD}_{\mathrm{diff}}(\mu,\nu;\mathcal{P})$ by computing the $\mathrm{MSD}_{\mathrm{diff}}$ on the empirical distributions, $\mathrm{MSD}_{\mathrm{diff}}(\hat{\mu},\hat{\nu};\mathcal{P})$.
@@
For any term $S\in\mathcal{S}_{\mathcal{P}}$ by the triangle inequality we have
@@
$\mathrm{MSD}_{\mathrm{diff}}(\mu,\nu;\mathcal{P})\leq$ $\mathrm{MSD}_{\mathrm{diff}}(\hat{\mu},\hat{\nu};\mathcal{P})$
$+\sup_{S\in\mathcal{S}_{\mathcal{P}}}\left|\hat{\mu}(S)-\mu(S)\right|$
$+\sup_{S\in\mathcal{S}_{\mathcal{P}}}\left|\hat{\nu}(S)-\nu(S)\right|.$
The quantity $\sup_{S\in\mathcal{S}_{\mathcal{P}}}\left|\hat{\mu}(S)-\mu(S)\right|$ describes the deviation of the empirical mean from the true mean over the $\mathcal{S}_{\mathcal{P}}$ family of functions. Therefore, by the uniform concentration results for bounded finite families (see, for instance, Theorem 2.13 in [37]), with probability at least $1-\delta$ over $\left\{x_{i}\right\}$, we have
$\sup_{S\in\mathcal{S}_{\mathcal{P}}}\left|\hat{\mu}(S)-\mu(S)\right|\leq 2\sqrt{\frac{\log\left|\mathcal{S}_{\mathcal{P}}\right|+\log\frac{2}{\delta}}{2N_{1}}}.$ (13)
@@
$\log\left|\mathcal{S}_{\mathcal{P}}\right|\leq 2\left|\mathcal{P}\right|.$ (14)
@@
3.2 MSD Estimation As Classification
To estimate the MSD for empirical distributions $\hat{\mu},\hat{\nu}$, it is more convenient to rephrase the maximization problem (5) as minimization of a classification loss, where the classifiers are taken from the family of terms, $\mathcal{S}_{\mathcal{P}}$. As we detail in Section 4, this point of view allows us to incorporate various useful ideas from the field of DNF classification into our MIO-based minimization algorithm.
@@ 4 Practical Estimation
Thus, there is a question of how to estimate the full $\mathrm{MSD}_{\mathrm{diff}}$ distance effectively. To this end, recall that subgroups $S\in\mathcal{S}_{\mathcal{P}}$ constitute terms, i.e., conjunctions of literals formed on a subset of features $\mathcal{P}$. And we also showed that one can formulate the $\mathrm{MSD}_{\mathrm{diff}}$ distance maximization using two classification problems.
Figure 1: Illustrative example. To naively look for a subgroup for which two distributions differ most, one would have to enumerate all subgroups. We show the number of subgroups in real datasets and the mean number of subgroups (subsets of the datasets). Each distance metric can be computed within 10-minute time frame. This is compared to the proposed MIO-based $\mathrm{MSD}_{\mathrm{diff}}$, which does not enumerate subgroups and thus scales better.
By using the methods of [51, 1, 53], we could learn a DNF that distinguishes between $\mu$ and $\nu$. Note that by definition, every term in such a DNF would correspond to some subgroup $S$ on which $\mu$ and $\nu$ differ. Since we search for a single member $S\in\mathcal{S}_{\mathcal{P}}$, we do not need to search for a full DNF, we rather need a single term – single conjunction.
To ensure that by minimizing the 0-1 loss, we indeed optimize the same notion as searching for the subgroup with maximal disparity, we must have balanced classes, i.e., an equal number of samples for each label. Alternatively, we can weigh the 0-1 losses of samples from a given class by $1/\text{number of samples of a given class}$. We can thus run any conjunction-learning algorithm to obtain a subgroup and the loss $L_{1}$, then flip the labels $y_{i}$, and optimize again to get the loss $L_{2}$. Then $\textrm{MSD}_{\text{diff}}(\hat{\mu},\hat{\nu};\mathcal{P})=1-\min(L_{1},L_{2})$, assuming the found solutions of the classification problems are globally optimal, i.e., the found subgroups (terms) have the minimal loss out of all in $\mathcal{S}_{\mathcal{P}}$.
4.1 MIO formulation
To find the globally optimal subgroup (i.e., conjunction or 1-term DNF), one can utilize Mixed-Integer Optimization (MIO). Let $\mathcal{D}_{+}$ and $\mathcal{D}_{-}$ be sets of indices of samples from $\mu$ and $\nu$, respectively, and let $\mathcal{D}=\mathcal{D}_{+}\cup\mathcal{D}_{-}$ be the set of indices of all samples. Our formulation is related to the 0-1 error DNF formulation of [45], but we search for a single term only and use a different objective, which leads to us also having to use further constraints.
@@ 4.1 MIO formulation
$\max\;$ $o$ (15a)
$\mathrm{s.t.}\;$ $o\leq\frac{1}{\left|\mathcal{D}_{+}\right|}\sum_{i\in\mathcal{D}_{+}}{\hat{y}_{i}}-\frac{1}{\left|\mathcal{D}_{-}\right|}\sum_{i\in\mathcal{D}_{-}}{\hat{y}_{i}}+2b$ (15b)
$o\leq\frac{1}{\left|\mathcal{D}_{-}\right|}\sum_{i\in\mathcal{D}_{-}}{\hat{y}_{i}}-\frac{1}{\left|\mathcal{D}_{+}\right|}\sum_{i\in\mathcal{D}_{+}}{\hat{y}_{i}}+2(1-b)$ (15c)
$\hat{y}_{i}\leq 1-(u_{j}-x_{i,j}\cdot u_{j})\hskip 28.45274pti\in\mathcal{D},\;j\in\mathcal{P}$ (15d)
$\hat{y}_{i}\geq 1-\sum_{j\in\mathcal{P}}(u_{j}-x_{i,j}u_{j})\hskip 51.21504pti\in\mathcal{D}$ (15e)
$\sum_{i\in\mathcal{D}}\hat{y}_{i}\geq N_{\text{min}}$ (15f)
$0\leq\hat{y}_{i}\leq 1\hskip 116.65646pti\in\mathcal{D}$ (15g)
$u_{j},b\in\{0,1\}\hskip 105.2751ptj\in\mathcal{P},$ (15h)
where $u_{j}$ is equal to 1 if and only if feature $j$ is present in the conjunction and $\hat{y}_{i}\in\{0,1\}$ are variables representing whether each sample belongs to the subgroup (i.e., is classified as 1). Indeed, $\hat{y}_{i}$ takes only binary values, due to the constraints on its value, together with the fact that $x_{i,j}$ and $u_{j}$ are binary.
The first two constraints—(15b) and (15c)—formulate the absolute value objective. The binary variable $b$ serves to relax one of the upper bounds, to allow the objective variable $o$ to be constrained by the higher bound out of the two. The value 2 is a tight “big-M” value in this context since the difference of the two means takes values from [1,1][-1,1].
@@
By solving this MIO problem, we find the subgroup with the highest absolute difference in probability of belonging to $\mu$ or $\nu$, i.e., the empirical estimate of $\textrm{MSD}_{\mathrm{diff}}(\mu,\nu;\mathcal{P})$.
5 Experiments
Our main goal is to measure the distance between distributions, focusing on protected subgroups. We thus wish to compare $\mathrm{MSD}_{\mathrm{diff}}$ to other measures of distances on distributions, as per Table 1. Since the data we consider is binary or categorical, we utilize the overlap kernel function $k(x_{i_{1}},x_{i_{2}})=\sum_{j=1}^{d}\left\llbracket x_{i_{1},j}=x_{i_{2},j}\right\rrbracket/d$ for the MMD evaluation. We use our own implementation of TV and MMD, and for W1 and W2, we utilize POT [15], an open-source Python library for the computation of optimal transport.
Additionally, we would like to compare the evaluation of $\mathrm{MSD}_{\mathrm{diff}}$ utilizing the proposed MIO formulation to other DNF optimization algorithms. We choose the classical algorithm Ripper [6] and BRCG [10] as a modern DNF learner. Since there is no public implementation of the original BRCG, we take the non-MIO re-implementation of BRCG, called BRCG-light [1] and the Ripper algorithm available in the AIX360 library [1]. We modify both implementations to return a single-term DNF, to ensure comparability to our method. As addressed earlier, we run each method two times, once to minimize error and the second time to maximize it (minimize the error with flipped labels). In case one of the distributions has more samples, we subsample it to have an equal number of samples for each distribution. This makes minimizing the 0-1 loss equivalent to finding the $\mathrm{MSD}_{\mathrm{diff}}$.
To solve the proposed MIO formulation (15), we model it using the Pyomo library [3] allowing easy use of a variety of solvers. We use the Gurobi solver [20]. We set the $N_{\text{min}}$ parameter to 10, allowing only subgroups with at least ten samples, although one could also use (13) and a fixed probability level $\delta$.
5.1 Datasests

### v2

@@ 2.4. Mixed-integer optimization
Mixed-integer optimization (MIO, (Wolsey, 2021)) is a powerful framework for modeling and solving mathematical optimization problems, where some decision variables take values from a discrete set while others are continuously valued. Despite MIO being a general framework for solving NP-hard problems, the MIO solvers are speeding up by approximately 22% every year, excluding hardware speedups (Koch et al., 2022). We use the abbreviation MIO, though we consider only mixed-integer linear formulations.
@@
MIO is used in machine learning, especially when one optimizes over discrete measures or decisions. This includes learning logical rules, including DNFs. Malioutov and Varshney (2013) learn DNFs through a sequential generation of terms using an LP relaxation of an MIO formulation. Later, Wang and Rudin (2015) used MIO to optimize full DNFs, and Su et al. (2016) introduced a formulation with Hamming distance as an alternative objective function. A crucial improvement to the scalability of exact learning of DNFs was the BRCG (Dash et al., 2018), which utilizes column generation to generate candidate terms. Recently, MIO was utilized to learn a DNF classifier with fairness constraints (Lawless et al., 2023). Importantly, we optimize only a single term (conjunction), representing the maximally discrepant subgroup, to evaluate bias. We do not perform predictions.
@@
As an aside, note that (Nair et al., 2021) (see also (Haldar et al., 2023)) uses DNFs to compare two models. This is done by building two DNFs as proxy models and comparing them. However, we could also directly compare the distributions of the two models, as proposed here.
3. Maximum Subgroup Discrepancy
@@
Let $\mathcal{X}\subset\mathbb{R}^{d}$ be an input space with $d$ features, and let $\mathcal{P}\subseteq[d]$ be a subset of features that are protected attributes. Throughout the rest of the paper, we assume that the protected attributes are binary, which can always be achieved by quantization and one-hot encodings. For a vector $x\in\mathbb{R}^{d}$ and $p\in\mathcal{P}$ let $x_{p}\in\left\{0,1\right\}$ denote its $p$-th coordinate, and let $\bar{x}_{p}:=1-x_{p}$ denote its logical negation. The set of functions $\mathcal{L}=\left\{x_{p},\bar{x}_{p}\right\}_{p\in\mathcal{P}}$ is called literals. A term is a conjunction of literals. That is, for a subset $S\subset\mathcal{L}$, the term $\chi_{S}$ is defined as a function $\mathbb{R}^{d}\rightarrow\left\{0,1\right\}$ given by
(3) $\chi_{S}(x)=\prod_{s\in S}s(x).$
@@
The subsets $S\subset\mathcal{L}$ are called subgroups, and with some abuse of notations, we also refer to the corresponding parts of the data defined by $S$,
(4) $\mathcal{X}_{S}=\left\{x\in\mathbb{R}^{d}\hskip 5.69054pt|\hskip 5.69054pt\chi_{S}(x)=1\right\},$
@@
With this notation, the $\mathrm{MSD}$ distance on feature set $\mathcal{P}$ between two distributions, $\mu,\nu$, is defined as
(5) $\mathrm{MSD}(\mu,\nu;\mathcal{P})=\sup_{S\in\mathcal{S}_{\mathcal{P}}}\left|\mu(\chi_{S})-\nu(\chi_{S})\right|.$
@@
(6) $\mathcal{B}_{\mathcal{P}}=\left\{\prod_{p\in\mathcal{P}}l_{p}\hskip 5.69054pt|\hskip 5.69054pt\text{ where $l_{p}=x_{p}$ or $l_{p}=\bar{x}_{p}$}\right\}\subset\mathcal{S}_{\mathcal{P}}.$
@@
(7) $\left\|\mu-\nu\right\|_{\infty}=\sup_{S\in\mathcal{B}}\left|\mu(\chi_{S})-\nu(\chi_{S})\right|.$
@@
Let $\mathcal{F}_{\mathcal{P},\infty}$ be the set of functions $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ that depend only on the coordinates in $\mathcal{P}$, and are bounded by $1$. Then, the total variation111marginalised to $\mathcal{P}$ (TV), may be written as
(8) $\left\|\mu-\nu\right\|_{TV}$ $=\frac{1}{2}\sup_{f\in\mathcal{F}_{\mathcal{P},\infty}}\left|\mu(f)-\nu(f)\right|$
(9) $=\frac{1}{2}\sum_{S\in\mathcal{B}}\left|\mu(\chi_{S})-\nu(\chi_{S})\right|,$
@@
(10) $\left\|\mu-\nu\right\|_{\infty}\leq\mathrm{MSD}(\mu,\nu;\mathcal{P})\leq\left\|\mu-\nu\right\|_{TV},$
@@
We thus observe that $\mathrm{MSD}(\mu,\nu;\mathcal{P})$ is a stronger distance than $\ell_{\infty}$, which allows us to consider subgroups with partially specified values. At the same time, it is not as strong as the total variation, which requires the sum over all base term differences to be small (rather than each of the differences being small individually). On the other hand, as discussed in Section 1, TV is, in fact, too strong to be practically computable due to its high sample complexity, while $\mathrm{MSD}$ has a sample complexity linear in the number of attributes.
3.1. Sample Complexity
In this section, we prove Theorem 3.1, which allows us to quantify the error of estimating $\mathrm{MSD}(\mu,\nu;\mathcal{P})$ from finite samples of these distributions.
@@ 3.1. Sample Complexity
(11) $\hat{\mu}=\frac{1}{N_{1}}\sum_{i}\delta_{x_{i}}\text{ and }\hat{\nu}=\frac{1}{N_{2}}\sum_{i}\delta_{x^{\prime}_{i}},$
@@
(12) $\mathrm{MSD}(\mu,\nu;\mathcal{P})\leq\mathrm{MSD}(\hat{\mu},\hat{\nu};\mathcal{P})+4\sqrt{\frac{2\left|\mathcal{P}\right|+\log\frac{2}{\delta}}{2N}}.$
In words, if the number of samples $N$ is of the order of the number of protected attributes, $\left|\mathcal{P}\right|$, or larger, we can estimate $\mathrm{MSD}(\mu,\nu;\mathcal{P})$ by computing the $\mathrm{MSD}$ on the empirical distributions, $\mathrm{MSD}(\hat{\mu},\hat{\nu};\mathcal{P})$.
@@
For any term $S\in\mathcal{S}_{\mathcal{P}}$, by the triangle inequality we have
@@
$\mathrm{MSD}(\mu,\nu;\mathcal{P})\leq\;$ $\mathrm{MSD}(\hat{\mu},\hat{\nu};\mathcal{P})$
$+\sup_{S\in\mathcal{S}_{\mathcal{P}}}\left|\hat{\mu}(S)-\mu(S)\right|+\sup_{S\in\mathcal{S}_{\mathcal{P}}}\left|\hat{\nu}(S)-\nu(S)\right|.$
The quantity $\sup_{S\in\mathcal{S}_{\mathcal{P}}}\left|\hat{\mu}(S)-\mu(S)\right|$ describes the deviation of the empirical mean from the true mean over the $\mathcal{S}_{\mathcal{P}}$ family of functions. Therefore, by the uniform concentration results for bounded finite families (see, for instance, Theorem 2.13 in (Mohri et al., 2018)), with probability at least $1-\delta$ over $\left\{x_{i}\right\}$, we have
(13) $\sup_{S\in\mathcal{S}_{\mathcal{P}}}\left|\hat{\mu}(S)-\mu(S)\right|\leq 2\sqrt{\frac{\log\left|\mathcal{S}_{\mathcal{P}}\right|+\log\frac{2}{\delta}}{2N_{1}}}.$
@@
(14) $\log\left|\mathcal{S}_{\mathcal{P}}\right|\leq 2\left|\mathcal{P}\right|.$
@@
3.2. MSD Estimation As Classification
To estimate the $\mathrm{MSD}$ for empirical distributions $\hat{\mu},\hat{\nu}$, it is more convenient to rephrase the maximization problem (5) as minimization of a classification loss, where the classifiers are taken from the family of terms, $\mathcal{S}_{\mathcal{P}}$. As we detail in Section 4, this point of view allows us to incorporate various useful ideas from the field of DNF classification into our MIO-based minimization algorithm.
@@ 4. Practical Estimation
Thus, there is a question of how to estimate the full $\mathrm{MSD}$ distance effectively. To this end, recall that subgroups $S\in\mathcal{S}_{\mathcal{P}}$ constitute terms, i.e., conjunctions of literals formed on a subset of features $\mathcal{P}$. We also showed that one can formulate the $\mathrm{MSD}$ distance maximization using two classification problems.
By using the methods of (Wang and Rudin, 2015; Arya et al., 2019; Wei et al., 2019), we could learn a DNF that distinguishes between $\mu$ and $\nu$. Note that by definition, every term in such a DNF would correspond to some subgroup $S$. Since we search for a single member $S\in\mathcal{S}_{\mathcal{P}}$, we do not need to search for a full DNF, we rather need a single term – single conjunction.
To ensure that by minimizing the 0-1 loss, we indeed optimize the same notion as searching for the subgroup with the maximal discrepancy, we must have balanced classes, i.e., an equal number of samples for each label. Alternatively, we can weigh the 0-1 losses of samples from a given class by $1/\text{number of samples of a given class}$. Thus, we can run any conjunction-learning algorithm to obtain a subgroup and the loss $L_{1}$, flip the labels $y_{i}$, and optimize again to get the loss $L_{2}$. Then $\textrm{MSD}(\hat{\mu},\hat{\nu};\mathcal{P})=1-\min(L_{1},L_{2})$, assuming the found solutions of the classification problems are globally optimal, i.e., the found subgroups (terms) have the minimal loss out of all in $\mathcal{S}_{\mathcal{P}}$.
4.1. MIO formulation
To find the globally optimal subgroup (i.e., conjunction or 1-term DNF), one can utilize Mixed-Integer Optimization (MIO). Let $\mathcal{D}_{+}$ and $\mathcal{D}_{-}$ be sets of indices of samples from $\mu$ and $\nu$, respectively, and let $\mathcal{D}=\mathcal{D}_{+}\cup\mathcal{D}_{-}$ be the set of indices of all samples. Our formulation is related to the 0-1 error DNF formulation of (Su et al., 2016), but we search for a single term only and use a different objective, which leads to us also having to use further constraints.
@@ 4.1. MIO formulation
(15a) $\max\;$ $o$
(15b) $\mathrm{s.t.}\;$ $o\leq\frac{1}{\left|\mathcal{D}_{+}\right|}\sum_{i\in\mathcal{D}_{+}}{\hat{y}_{i}}-\frac{1}{\left|\mathcal{D}_{-}\right|}\sum_{i\in\mathcal{D}_{-}}{\hat{y}_{i}}+2b$
(15c) $o\leq\frac{1}{\left|\mathcal{D}_{-}\right|}\sum_{i\in\mathcal{D}_{-}}{\hat{y}_{i}}-\frac{1}{\left|\mathcal{D}_{+}\right|}\sum_{i\in\mathcal{D}_{+}}{\hat{y}_{i}}+2(1-b)$
(15d) $\hat{y}_{i}\leq 1-(u_{j}-x_{i,j}\cdot u_{j})\hskip 28.45274pti\in\mathcal{D},\;j\in\mathcal{P}$
(15e) $\hat{y}_{i}\geq 1-\sum_{j\in\mathcal{P}}(u_{j}-x_{i,j}u_{j})\hskip 51.21504pti\in\mathcal{D}$
(15f) $\sum_{i\in\mathcal{D}}\hat{y}_{i}\geq N_{\text{min}}$
(15g) $0\leq\hat{y}_{i}\leq 1\hskip 116.65646pti\in\mathcal{D}$
(15h) $u_{j},b\in\{0,1\}\hskip 105.2751ptj\in\mathcal{P},$
where $u_{j}$ is equal to 1 if and only if feature $j$ is present in the conjunction, and $\hat{y}_{i}\in\{0,1\}$ are variables representing whether each sample belongs to the subgroup (i.e., is classified as 1). Indeed, $\hat{y}_{i}$ takes only binary values, due to the constraints on its value, together with the fact that $x_{i,j}$ and $u_{j}$ are binary.
The first two constraints—(15b) and (15c)—formulate the absolute value objective. The binary variable $b$ serves to relax one of the upper bounds to allow the objective variable $o$ to be constrained by the higher bound out of the two. The value 2 is a tight “big-M” value in this context since the difference of the two means takes values from [1,1][-1,1].
@@
By solving this MIO problem, we find the subgroup with the highest absolute difference in probability of belonging to $\mu$ or $\nu$, i.e., the empirical estimate of $\textrm{MSD}(\mu,\nu;\mathcal{P})$.
5. Experiments
Experimentally, we aim to answer the following questions:
  • Q1
    Can the use of mixed-integer programming in (15) be efficient? What is the run time in practice?
  • Q2
    Does $\mathrm{MSD}$ have a low sample complexity, as predicted by Theorem 3.1?
  • Q3
    Could one use alternative algorithms for learning DNF in estimating $\mathrm{MSD}$?
Notice that the answers to all three questions are non-obvious. In Q1, we wish to compare different algorithms for evaluating MSD, whose run-time is exponential in the number of protected attributes.
In Q2, we want to corroborate our sample complexity results of Section 3.1. We thus wish to compare $\mathrm{MSD}$ to other measures of distances on distributions, as per Table 1.
In Q3, we would like to compare the evaluation of $\mathrm{MSD}$ utilizing the proposed MIO formulation to other DNF learning algorithms, even if not utilized in bias detection previously. We choose the classical algorithm Ripper (Cohen, 1995) and the more modern BRCG (Dash et al., 2018).
5.1. Datasests